OpenStudy (anonymous):

As a first-time homeowner, there are going to many decisions that you need to make, such as whom to hire for the upkeep of your lawn. You have just received advertisements from 2 companies in your area: Greener Lawns and Lawns for Less. Greener Lawns charges an initial fee of $200, plus $50 per hour for labor costs. Lawns for Less charges an initial fee of $300, plus $25 per hour for labor costs. 1. If h represents the number of labor hours and C represents the cost, write the total cost equation for Greener Lawns. 2. If h represents the number of labor hours and C represents the cost, write t

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

f(h)=50h + 200 or 50h +C, where C=200 Cannot see the last bit, is there a 2nd question?

OpenStudy (anonymous):

As a first-time homeowner, there are going to many decisions that you need to make, such as whom to hire for the upkeep of your lawn. You have just received advertisements from 2 companies in your area: Greener Lawns and Lawns for Less. Greener Lawns charges an initial fee of $200, plus $50 per hour for labor costs. Lawns for Less charges an initial fee of $300, plus $25 per hour for labor costs. 1. If h represents the number of labor hours and C represents the cost, write the total cost equation for Greener Lawns. 2. If h represents the number of labor hours and C represents the cost, write the total cost equation for Lawns for Less. 3. Solve the system of equations for the total cost of lawn care using the desired technique: substitution, elimination, or graphing. 4. Document how you came to these conclusions for accuracy.

OpenStudy (anonymous):

1. C=50h+200 2. C=25h +300 3. elimination is the easiest. subtract the 2nd from the first: C-C=50h-25h-100 25h=100 h=4 put this back to 1. or 2. to get C=400

OpenStudy (anonymous):

To write it with words, you can say that if the work is for 4 hours than the cost is equal (400) and it does not matter which company you choose